solve each system using substitution.

4x=3y-2
18=3x+y

Answer by TutorDelphia(193)

First we need to find out what x equals or y equals...it looks like it will be easy to get y alone in the second equation:
4x=3y-2
18=3x+y
18-3x=y
Now that we know y equals 18-3x I can plug in 18-3x anywhere I see a y
4x=3(18-3x)-2
Distribute the 3
4x=54-9x-2
combine like terms
4x=52-9x
Add 9x to both sides
13x=52
divide both sides by 13
x=4
Now that I know x=4 I can plug it into one of the equations:
4*4=3y-2
16=3y-2
Add 2 to both sides
18=3y
divide both sides by 3
y=6